a) Why is a monopolistically competitive firm less efficient than a perfectly competitive firm?
It produces at an output that is lower than its minimum efficient scale (MES)
It earns positive economic profits in the long run
It deters entry of new firms by putting up entry barriers
All of the answers are correct
b) Suppose a monopolistically competitive firm has MC=4Q+5. Its demand is P=145-3Q and marginal revenue is MR=145-6Q. What is its profit-maximizing output level?
17
14
16
15
c) Continue with part (b): What is its profit?
-$650
$980
-$525
$775
d) Game theory indicates
It is not optimal for a firm to cheat on the other
That competitive behavior of firms leads to higher joint profits for the firms in comparison to the cooperative outcome
Collusion is a self-enforcing Nash equilibrium
That competitive behavior of firms leads to lower joint profits for the firms in comparison to the cooperative outcome
e) Suppose there are two firms in the market, Firm A and Firm B. The total market demand is given by P=40-2.5Q, MR=40-5Q, and MC=ATC=$10. What is the profit maximizing price for a cartel?
$20
$30
$35
$25
f) Continue with part (e): What is the output level for the cartel?
12
24
6
16
g) Continue with part (e): What is the profit for the cartel?
$80
$100
$110
$90
h) Which of the following statements best describe(s) an imperfectly competitive firm?
It is a price-setter
It faces a horizontal demand curve
It faces a fixed cost when producing its products in the long run
All of the answers are correct
i) The monopoly portion of the market structure "monopolistic competition" refers to _____, while the competition portion refers to _____.
Differentiated products; barriers to entry
Differentiated products; free entry and exit
Identical products; free entry and exit
Identical products; barriers to entry
j) Continue with part (i): What does the monopoly portion refer to?
Firms in the long run will earn zero economic profits
The demand curve that a typical firm faces is negatively sloped
All of the answers are correct
Because of brand loyalty, a firm can raise the price of its product without worrying about any of its customers will switch to buy other similar brands
Answer #1

A) Ans: All of the answer are correct.

In the long run, a firm in monopolistic competition produces less than the efficient scale and has excess capacity because the firm faces a downward sloping demand curve. Whether the Perfect competition always faces horizontal demand curve.

A monopolistic firm earn economic profit in long run but firm in perfect competition always earn normal profit in the long run, which shows less competition in monopolistic firms.

Barriers to entry is another factor which shows less competition in monopolistic firm.

B) Ans: 14

At Profit Maximizing Level, MR=MC

145 - 6Q = 4Q + 5

= Q = 14

So ans is 14

C) Ans : 980

Profit = TR - TC

TR = P × Q

= (145 - 3Q)Q

= 145Q - 3Q2

TC = \int (MC) dQ

= \int (4Q+5) dQ

= 2Q2 + 5Q

Profit = 145Q- 3Q2 - (2Q2+5Q)

Put Q = 14, we get

Profit = 980

D) Ans: Game theory indicates, the competative behavior of firms leads to lower joint profits for the firm in comparison to the cooperative outcome.

It is because Each firm knows that if they work together and do not cheat each other, they will be able to restrict output and raise prices, thereby enjoying above-normal profits.

E) Ans : $25

MR = MC

40-5Q = 10

Q = 6

Put Q in P = 40-2.5(6)

= P= 25

F) Ans: Output is Q = 6

As shown in part E.

G) Ans : Profit = TR - TC

TR = P × Q

= (40-2.5Q)Q = 40Q - 2.5Q2

Put Q = 6,

TR = 150

TC = ATC× Q

= 10×6

= 60

Profit = TR - TC

Profit = 150- 60 = 90

H) Ans: Imperfect Competitive firm is a Price-setter

It is because all imperfect competitive firm has some power of price setting.

Also imperfect competitive firms faces downward sloping demand curve not horizontal demand curve and also they does not face fixed cost in long run.

I) Ans : Differentiate products; Barriers to entry as monopolistic competition refers to differentiate products while competition portion refers to barriers to entry.

J) Ans : Monopoly portion refers to

Because of brand loyalty, a firm can raise the price of its product without worrying about any of its customers will switch to buy other similar brands.